OpenStudy (anonymous):

Find the x-, y-, and z-intercepts for the equation. Enter your answers separated by commas; example: -2, 3, 4. 9x + 3y - 6z = 18

OpenStudy (amistre64):

divide it all by 18 and simplify

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

k

OpenStudy (amistre64):

[9x + 3y - 6z = 18]/18 x/2 + y/6 - z/3 = 1 2 , 6 , -3

OpenStudy (anonymous):

do you always divide by the end number?

OpenStudy (amistre64):

it might be easier to see the negatives if you attach them to the bottoms like this: x y z --- + --- + --- = 1 2 6 -3

OpenStudy (amistre64):

It is one method in finding out the intercepts of an equation.

OpenStudy (amistre64):

a more popular method is simply to zero out the other variable and solve for the one thats left

OpenStudy (anonymous):

the one with zero u add it to the space where the variable is?

OpenStudy (amistre64):

yes, for example, to find the z intercept: 9(0) + 3(0) - 6z = 18 0+0 -6z = 18 -6z = 18 z = 18/-6 = -3

OpenStudy (anonymous):

can you write a example problem down and I try to solve it and you tell me if it's right?

OpenStudy (amistre64):

4x -5y -10z = 20

OpenStudy (anonymous):

5, -4, -2 is that right?

OpenStudy (amistre64):

that is correct :)

OpenStudy (anonymous):

wow thank you I was stuck on this for like a day

OpenStudy (amistre64):

youre welcome :)